# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

### v2

@@ 26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})\leftarrow\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\cdot\mathcal{N}(0,\mathbf{I})\right)$;
@@
5.1. Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al., 2016; Fu et al., 2023; Bu et al., 2022; Yang et al., 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (9), into Equation (12):
(12) $\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$
@@ 5.1. Fair-clipping Strategy for Gradient
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\lambda\cdot\Delta_{i}^{j}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$
(13) $=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$
@@
To achieve differential privacy protection, we replace the gradient of loss function by Equation (5.1) based on Equations (4)-(6), and reorganize the gradient descent equation for client $i$:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})}{\max\left(1,\frac{\left(1+\lambda\cdot\Delta_{i}^{j}\right)\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}{C}\right)}$
$\quad+\sigma C\cdot\mathcal{N}(0,\mathbf{I})]$
(14) $=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$
@@
(15) $C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$
Hence, our strategy can be regarded as an fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (15) corresponds to the traditional DP clipping method, and FedFDP no longer ensures fairness. For large $C$, Equation (15) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (15) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. Our conclusions can be verified from the experiments (Figure 4).
5.2. Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 2-2 of Algorithm 2, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
@@ 5.2. Adaptive Clipping Method for Loss
(16) $C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$
@@
(17) $\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$
@@ 5.3. Overall Algorithm
  • DPSGD (lines 2-2): In lines 2-2, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 2 and execute the gradient decent.
@@
  • DP for local loss (lines 2-2): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6. Optimal Fairness Parameter
@@
In this section, we examine the convergence of the FedFDP algorithm to identify an optimal fairness parameter $\lambda$ that minimizes the convergence bound.
6.1. Convergence Analysis
Assumption 1.
Assumption 2.
Assumption 3.
Theorem 1.
(18) $\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$
where:
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
The meaning of $\Gamma$ is consistent with (Li et al., 2020a), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive (18). For the detailed proof, please refer to Appendix A.1333Due to the lengthy proof, please refer to https://anonymous.4open.science/r/FedFDP-5607/APPENDIX/FedFDP_appendix.pdf for the appendix..
6.2. Optimal $\lambda$
Theorem 2.
Proof. When $t\rightarrow T$ and $T$ is sufficiently large, $\|\nabla F_{i}(\mathbf{w}_{t}^{i},x_{j})\|\rightarrow 0$, $C_{t}^{i,j}=1+\lambda\Delta_{i}^{j}$. Assuming $Q_{0}\leq\Delta_{i}^{j}\leq Q_{1}$, then $1+Q_{0}\lambda\leq C_{t}^{i,j}\leq 1+Q_{1}\lambda$. We rearrange the convergence upper bound in Equation (18) to obtain a function in terms of $\lambda$,
(19) $P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$
$\mathcal{F}^{\prime}(\lambda)$ $=\frac{(3a_{1}\lambda^{2}+2a_{2}\lambda+a_{3})(a_{5}\lambda+1)}{(a_{5}\lambda+1)^{2}}-\frac{a_{5}(a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4})}{(a_{5}\lambda+1)^{2}}$
(20) $=\frac{2a_{1}a_{5}\lambda^{3}+(a_{2}a_{5}+3a_{1})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}}{(a_{5}\lambda+1)^{2}}.$
Let the denominator of $\mathcal{F}^{\prime}(\lambda)$ be a new function:
(21) $\mathcal{G}(\lambda)=2a_{1}a_{5}\lambda^{3}+(3a_{1}+a_{2}a_{5})\lambda^{2}+2a_{2}\lambda+a_{3}-a_{4}a_{5}.$
$\Delta$ $=(3a_{1}+a_{2}a_{5})^{2}-3(2a_{1}a_{5}\cdot 2a_{2})$
$=9a_{1}^{2}+a_{2}^{2}a_{5}^{2}+6a_{1}a_{2}a_{5}-12a_{1}a_{2}a_{5}$
(22) $=(3a_{1}-a_{2}a_{5})^{2}\geq 0.$
The derivative of $\mathcal{G}(\lambda)$ is given by
(23) $\mathcal{G}^{\prime}(\lambda)=6a_{1}a_{5}\lambda^{2}+2(a_{2}a_{5}+3a_{1})\lambda+2a_{2}.$
It is evident that when $\lambda>0$, $\mathcal{G}^{\prime}(\lambda)>0$.
(24) $\mathcal{G}(x)=x^{3}+a_{6}x+a_{7},$
where
  • $a_{6}=\frac{-(3a_{1}-a_{2}a_{5})^{2}}{12a_{1}^{2}a_{5}^{2}}$,
[…]
7. Privacy Analysis
As Algorithm 2 shown, the $i$-th client will return model parameters $\mathbf{w}_{t+1}^{i}$ and loss value $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ before the server aggregates the model parameters in round $t+1$. $\mathbf{w}_{t+1}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ both access the private training data of local client, so they are both to perform differential privacy preservation, as in lines 2-2 and lines 2-2 of the Algorithm 2, respectively. We will analyze their privacy with RDP separately and finally combine their privacy loss to get the overall privacy loss of FedFDP algorithm.
7.1. Privacy Loss of $\mathbf{w}_{t+1}^{i}$
Theorem 1.
(25) $R_{model}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma^{2}}\right),$
where $\sigma$ is noise multiplier of the $\mathbf{w}_{t+1}^{i}$, and $\alpha>1$ is the order.
The detail of proof please refer Appendix A.2.
7.2. Privacy Loss of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$
Theorem 2.
(26) $R_{loss}^{i}(\alpha)=\frac{T}{\alpha-1}\sum_{k=0}^{\alpha}\left(\begin{array}[]{l}\alpha\\ k\end{array}\right)(1-q)^{\alpha-i}q^{k}\exp\left(\frac{k^{2}-k}{2\sigma_{l}^{2}}\right),$
The proof is similar to that of the $\mathbf{w}_{t+1}^{i}$, so we omit it.
7.3. Privacy Loss of FedFDP
Since both $\mathbf{w}_{T}^{i}$ and $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ access the training set, we need to combine their RDP sequentially using Definition 3, and then use Lemma 4 to convert it to $(\epsilon,\delta)$-DP. Lastly, we can get the Privacy loss of FedFDP as follows.
Theorem 3.
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 1, and $R_{loss}^{i}(\alpha)$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 2.
Lemma 0.
(Conversion from RDP to DP (Balle et al., 2020)). if a randomized mechanism $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ satisfies $(\alpha,R)$-RDP , then it satisfies$(R+\ln((\alpha-1)/\alpha)-(\ln\delta+\ln\alpha)/(\alpha-1),\delta)$-DP for any $10<\delta<$ 1.
8. Experiments
Table 2. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$).
Datasets MNIST FashionMNIST CIFAR10
Accuracy Fairness Accuracy Fairness Accuracy Fairness
Table 3. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
@@ 8. Experiments
Table 4. FedFair and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability.
@@
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
Table 5. FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
Herterogeneity Scalability
Datasets MNIST FashionMNIST CIFAR10
Method $Dir(0.1)$ $Dir(0.5)$ $Dir(1)$ 10 clients 20 clients 50 clients
@@
[…]
@@
This section outlines our experimental framework, beginning with the presentation of baselines and default parameter configurations. We then detail the experimental assessment of FedFair and FedFDP across dimensions of performance, fairness, heterogeneity, and scalability. Ablation studies focusing on the hyperparameters $\lambda$, $C$ and $\epsilon$ of FedFDP are also included. All tasks will be repeated five times using different random seeds, and the mean and standard deviation will be reported.
8.1. Default Settings
We conducted experiments on three computer vision (CV) datasets: MNIST (LeCun et al., 1998), FashionMNIST (Xiao et al., 2017), and CIFAR10 (Krizhevsky, 2009). And we comparised our algorithms to six baselines:
  1. (1)
    FedAvg(McMahan et al., 2017): FedAvg is the most classic algorithm in federated learning. In fact, we can set $\lambda$ to 0, transforming FedFair and FedFDP into the FedAvg algorithm.
  2. (2)
    SCAFFLOD(Karimireddy et al., 2020): SCAFFOLD uses control variates (variance reduction) to correct for the “client-drift” in its local updates.
  3. (3)
    FedProx(Li et al., 2020b): FedProx introduces a proximal term to the objective function to address the issues of systems heterogeneity and statistical heterogeneity in federated learning.
  4. (4)
    FedDyn(Acar et al., 2021): FedDyn adds linear and quadratic penalty terms to the objective function to solve the problem of inconsistency between the minima of the local-device level empirical loss and those of the global empirical loss.
  5. (5)
    ALI-DPFL(Ling et al., 2023): ALI-DPFL dynamically obtains the optimal number of local DPSGD iterations for clients between any two sequential global updates through minimizing the convergence bound.
  6. (6)
    q-FFL(Li et al., 2020c): q-FFL amplifies loss values for clients with poorer performance by assigning them larger update weights, aiming to improve their model performance and achieve a balanced performance across multiple devices.
For all datasets, we adopted a widely used heterogeneous settings (Lin et al., 2020; Li et al., 2021a) to 10 clients which controlled by a Dirichlet distribution denoted as $Dir(\beta)$, and the default value of $\beta=0.1$ (Lin et al., 2020; Wang et al., 2020). We use a 4-layer CNN architecture (McMahan et al., 2017) which consists of two convolutional layers and two fully connected layers as model architecture.
8.2. Results
In this subsection, we systematically evaluate the performance and fairness of the FedFair and FedFDP under varying heterogeneity levels, expanding from $Dir(0.1)$ to both $Dir(0.5)$ and $Dir(1)$. We also assess their scalability by scaling the number of clients from the standard 10 to encompass 20 and 50 clients. Furthermore, we examine the influence of three pivotal hyperparameters within the FedFDP algorithm, namely, the adjustments to $\lambda$, $C$ and $\epsilon$.
8.2.1. FedFair
Performance and Fairness. Table 2 presents the experimental results of the FedFair algorithm compared to baseline methods. FedFair significantly enhances fairness while maintaining substantial model performance across the MNIST, FashionMNIST, and CIFAR10 datasets, with respective increases of 17.0%, 25.7%, and 9.6%.
Heterogeneity. Table 4 shows that FedFair outperforms baseline methods in accuracy while maintaining significant fairness when faced with diverse data distributions exhibiting varying levels of heterogeneity. Except for the $Dir(0.5)$ scenario, where it falls slightly short of q-FFL, FedFair demonstrated fairness improvements of 22.0% and 50% over the best baseline in $Dir(0.1)$ and $Dir(1.0)$, respectively.
8.2.2. FedFDP

### v3

@@ 26            $f_{i,j}=\min({C_{l}^{i,t}},\max(0,F_{i}(\mathbf{w}_{t+1}^{i},\xi_{j})))$;
28      $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}f_{i,j}+\sigma_{l}C_{l}^{i,t}\mathcal{N}(0,\mathbf{I})\right)$;
@@
5.1 Fair-clipping Strategy for Gradient
In the general DPSGD algorithm (Abadi et al. 2016; Fu et al. 2023; Bu et al. 2022; Yang et al. 2022), per-sample clipping is an essential process for obtaining sensitivity. In order to be compatible with per-sample clipping, initially, we expand our objective function of client $i$ in FedFair, Equation (5), into Equation (8):
$\min_{\mathbf{w}_{t}^{i}}H_{i}(\mathbf{w}_{t}^{i},\xi_{j})\triangleq F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}_{t}^{i},\xi_{j})-F(\mathbf{w}_{t}^{i})\right)^{2},$ (8)
@@ 5.1 Fair-clipping Strategy for Gradient
$\nabla H_{i}(\mathbf{w}_{t}^{i},\xi_{j})=\left(1+\lambda\cdot\Delta_{i}^{j}\right)\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j}),$ (9)
@@
To ensure the algorithm is differentially private, we replace the gradient of loss function by Equation (9), and reformulate the computation of the gradient descent for client $i$ as follows:
$\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\frac{\eta}{|\mathcal{B}_{i}|}[\sum^{|\mathcal{B}_{i}|}_{j=1}C^{i,j}_{t}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})+\sigma C\cdot\mathcal{N}(0,\mathbf{I})],$ (10)
@@
$C^{i,j}_{t}=\min\left(1+\lambda\cdot\Delta_{i}^{j},\frac{C}{\|\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})\|}\right).$ (11)
Hence, our strategy can be considered as a fair-clipping technique, tailored to meet the demands of differential privacy while also nurturing fairness within the model. When $\lambda=0$, Equation (11) corresponds to the traditional DP clipping method without fairness. For large $C$, Equation (11) simplifies to the first term, prompting FedFDP to adjust towards enhanced fairness. Conversely, for small $C$, Equation (11) simplifies to the second term, resulting in the gradient being clipped to the bound $C$. The impact of $C$ on the performance of FedFDP will be further discussed in the experiments.
5.2 Adaptive Clipping Method for Loss
Since the calculation of $\Delta_{i}^{j}$ requires clients to upload $F_{i}(\mathbf{w}_{t}^{i})$, we need to incorporate differential privacy to ensure that the algorithm does not compromise privacy. As shown in lines 1-1 of Algorithm 1, We need to clip $F_{i}(\mathbf{w}_{t}^{i})$ to the interval $[0,C_{l}^{i,t}]$, and then add Gaussian noise with a mean of 00 and a standard deviation of $\sigma_{l}\cdot C_{l}^{i,t}$ to ensure differential privacy. However, setting an appropriate clipping bound $C_{l}^{i,t}$ for $F_{i}(\mathbf{w}_{t}^{i})$ is challenging. A large $C_{l}^{i,t}$ avoids loss of fidelity to the original values, but it also leads to large injected noise.
@@ 5.2 Adaptive Clipping Method for Loss
$C_{l}^{i,t}=\frac{\sum_{j=1}^{|\mathcal{B}_{i}|}\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)+\mathcal{N}\left(0,({C_{l}^{i,t-1}}\cdot{\sigma_{l}})^{2}\right)}{|\mathcal{B}_{i}|},$ (12)
@@
$\operatorname{clip}\left(F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j})\right)=\min({C_{l}^{i,t-1}},\max(0,F_{i}(\mathbf{w}_{t-1}^{i},\xi_{j}))).$ (13)
@@ 5.3 Overall Algorithm
  • DPSGD (lines 1-1): In lines 1-1, we compute the fair-clipping bound $C_{t}^{i,j}$ and the per-sample gradient $\nabla F_{i}(\mathbf{w}_{t}^{i},\xi_{j})$ to get the processed gradient $\tilde{g}_{t}^{i,j}$. Then, we add mean 0 noise to the sum of per-sample processed gradient $\tilde{g}_{t}^{i,j}$ in line 1 and execute the gradient decent.
@@
  • DP for local loss (lines 1-1): As described in 5.2, add noise to the clipped local loss by using adaptive clipping bound $C_{l}^{i,t}$.
6 Finding Optimal Fairness Parameter
@@
In this section, we discuss how to identify an optimal fairness parameter $\lambda$ through the convergence analysis of the FedFDP algorithm.
@@
Table 1: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$), while the privacy budget $\epsilon=3.52$.
6.1 Convergence Analysis
Assumption 1
Assumption 2
Assumption 3
Theorem 1
$\mathbb{E}[F(\mathbf{w}_{t})]-F^{*}\leq\frac{L}{2t}\left(\frac{A}{\mu^{2}(2C_{t}-1)}+\mathbb{E}\|\mathbf{w}_{1}-\mathbf{w}^{*}\|^{2}\right),$ (14)
where:
  • $\hat{B}\triangleq\min_{i}|\mathcal{B}_{i}|$,
  • $\Gamma\triangleq F^{*}-\sum_{i=1}^{N}p_{i}F_{i}^{*}$,
The meaning of $\Gamma$ is consistent with (Li et al. 2020c), where a larger value of $\Gamma$ indicates that the data among different clients is more Non-IID.
proof sketch: We investigate the relationship between $\mathbb{E}\|\mathbf{w}_{t+1}-\mathbf{w}^{*}\|^{2}$ and $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$, and then use mathematical induction to obtain an upper bound for $\mathbb{E}\|\mathbf{w}_{t}-\mathbf{w}^{*}\|^{2}$. Finally, by utilizing Assumption 1, we derive Equation (14). For the detailed proof, please refer to Appendix.
@@ 6.1 Convergence Analysis
Method $\mathrm{Dir}(0.1)$ $\mathrm{Dir}(0.5)$ $\mathrm{Dir}(1)$ 10 clients 20 clients 50 clients
@@
Table 2: FedFDP and baselines: results of test accuracy (%) and fairness ($\Psi$) in different heterogeneity and scalability, while the privacy budget $\epsilon=3.52$.
@@
[…]
6.2 Optimal Fairness Parameter $\lambda$
Based on the convergence analysis, we find the “best” fairness parameter $\lambda$ that can minimize the convergence bound. We have the following theorem to do so:
Theorem 2
proof sketch: By reorganizing the convergence upper bound of Equation (14) into a function of $\lambda$:
$P(\lambda)=\frac{L}{2\mu T}\frac{a_{1}\lambda^{3}+a_{2}\lambda^{2}+a_{3}\lambda+a_{4}}{a_{5}\lambda+1},$ (15)
By using numerical analysis methods, the minimum value $P(\lambda)$ can be solved. For the meaning of $a_{1,2,3,4,5}$ and the proof process, please refer to the Appendix.
7 Privacy Analysis
Because each client has the same privacy loss in every round. As a result, we only assess each client’s privacy loss from the perspective of a single client. Following that, we examine the privacy loss of FedPDF for the $i$-th client as follows.
Theorem 3
[…]
where $10<\delta<$ 1, $R_{model}^{i}(\alpha)$ is the RDP of $w_{t+1}^{i}$ is computed by Theorem 4 in Appendix, and $R_{loss}^{i}(\alpha$ is the RDP of $\tilde{F}_{i}(\mathbf{w}_{t+1}^{i})$ which is computed by Theorem 5 in Appendix.
The details of the proof please refer to Appendix.
8 Experiments
This section outlines our experimental framework. We detail the experimental assessment of FedFDP across dimensions of performance, fairness, heterogeneity, and scalability, including ablation experiments focusing on the fairness hyperparameter $\lambda$, privacy budget $\epsilon$, clipping bound $C$ and noise multiplier $\sigma$.
In the appendix, we also provide an experimental assessment of FedFair across the same dimensions of performance, fairness, heterogeneity, and scalability. For an introduction to baselines, default values of hyperparameters, and the running environment, please refer to the Appendix.
All tasks were repeated five times and report the mean and standard deviation.
8.1 Results